OpenStudy (anonymous):

"hang time" of a punt is measured to be 4.50 seconds if he ball was kicked at an angle of 63 degrees above the horizontal and its caught at the same level which it was kicked, what was its initial speed?

OpenStudy (john_es):

Using kinematic equations, \[y=y_0+v_0\sin\alpha t-\frac{1}{2}gt^2\]And using \[y=y_0=0; t=4.50\ s;\ \ g=9.8 \ m/s^2\] \[v_0=\frac{gt}{2\sin\alpha}=24.7\ m/s\]